# Inclusions of real rank zero

### v2

@@ 2. Inclusions of commutative $\mathrm{C}^{*}$-algebras
The if direction follows from Lemma 1.11 and [10, Proposition 1.1]. Conversely, let us suppose that the inclusion has real rank zero. We first claim that $\phi$ is constant on each connected component of $Y$. To prove it, let $Y_{0}$ be a connected component of $Y$ and $y_{1},y_{2}\in Y_{0}$. If $Y_{0}$ is a singleton, we are done, so let us assume that $y_{1}$ and $y_{2}$ are distinct. Suppose for a contradiction that $\phi(y_{1})\neq\phi(y_{2})$. By Urysohn’s lemma, we then pick a self-adjoint $f\in C(X)$ such that $f(\phi(y_{1}))<0$ and $f(\phi(y_{2}))>0$. Let $\epsilon>0$ such that $2\epsilon<\min\{f(\phi(y_{2})),|f(\phi(y_{1}))|\}$.
Since $\phi^{*}$ has real rank zero, there exists an invertible self-adjoint $g\in C(Y)$ such that $\|g(y)-\phi^{*}(f)(y)\|\leq\epsilon$ for all $y\in Y$. In particular,
@@
commutes, where $\psi^{*}$ and $\pi^{*}$ are the dual maps of $\psi$ and $\pi$ respectively. Since $Y$ is compact and Hausdorff, it follows that $Y_{c}$ is compact and totally disconnected. It is well-known that $Y_{c}$ is Hausdorff. To see this, note that in any compact Hausdorff space, connected components are quasicomponents and hence closed (see for example [18, Proposition 1.4]). As $Y$ is compact and Hausdorff, and in particular normal, any two connected components are contained in disjoint open sets. Consequently, the quotient space $Y_{c}$ is Hausdorff. Hence, $C(Y_{c})$ has real rank zero by [10, Proposition 1.1]. Moreover, $\pi$ is surjective, which implies that $\pi^{*}$ is injective by duality. Hence $\pi^{*}(C(Y_{c}))$ has real rank zero. Since $C(X)\subseteq\pi^{*}(C(Y_{c}))\subseteq C(Y)$, we reach the conclusion. ∎
@@
Let $X,Y$ be compact Hausdorff spaces and $\phi:Y\to X$ be a continuous map which induces a $*{}^{*}$-homomorphism $\phi^{*}:C(X)\to C(Y)$. Then $\phi^{*}$ has real rank zero if and only if it factors through a commutative $\mathrm{C}^{*}$-algebra of real rank zero.
@@
The if direction follows from Lemma 1.10. Conversely, suppose that $\phi^{*}$ has real rank zero. Note that $\phi^{*}(C(X))$ is a commutative $\mathrm{C}^{*}$-algebra, so we can identify it with some $\mathrm{C}^{*}$-algebra $C(Z)$ by Gelfand duality. Then the conclusion follows by applying Theorem 2.1 to the inclusion $C(Z)\subseteq C(Y)$ and [10, Proposition 1.1]. ∎
We use Corollary 2.2 to characterise $*{}^{*}$-homomorphisms between commutative $\mathrm{C}^{*}$-algebras which have nuclear dimension equal to zero. Let us first record the following definition.
@@
([49, 8]) Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras and $\theta:A\to B$ be a $*{}^{*}$-homomorphism. It is said that $\theta$ has nuclear dimension equal to zero if for any finite set $\mathcal{F}\subseteq A$ and $\epsilon>0$, there exist a finite dimensional $\mathrm{C}^{*}$-algebra $F$ and cpc maps
@@
Moreover, a $\mathrm{C}^{*}$-algebra $A$ is said to have nuclear dimension equal to zero if the identity map on $A$ has nuclear dimension equal to zero. It is known that a separable $\mathrm{C}^{*}$-algebra has nuclear dimension equal to zero if and only if it is approximately finite dimensional ([51, Theorem 4.2.3]).
@@
Let $A,B$ be unital $\mathrm{C}^{*}$-algebras and $\theta:A\to B$ be a unital $*{}^{*}$-homomorphism with nuclear dimension equal to zero. Then $\theta$ has real rank zero.
@@
Let $a\in A$ be a self-adjoint contraction and $\epsilon>0$. Building on [51, Theorem 3.4], there exist a finite dimensional $\mathrm{C}^{*}$-algebra $F$, a ucp map $\psi:A\to F$, and a unital $*{}^{*}$-homomorphism $\eta:F\to B$ such that $\|\theta(a)-\eta(\psi(a))\|<\epsilon/2$ ([13, Lemma 3.1]). [^6] Even though [13, Lemma 3.1] is only stated in the setting when $A$ and $B$ are separable, this assumption is not needed. Since $\psi(a)$ is self-adjoint and $F$ is finite dimensional, there exists a self-adjoint invertible $b^{\prime}\in F$ such that $\|\psi(a)-b^{\prime}\|\leq\epsilon/2$. Therefore, $\eta(b^{\prime})$ is self-adjoint invertible and
@@
Let $X,Y$ be compact Hausdorff spaces and $\phi:Y\to X$ be a continuous map which induces a $*{}^{*}$-homomorphism $\phi^{*}:C(X)\to C(Y)$. The following are equivalent:
@@
The fact that (i) implies (ii) follows from Proposition 2.4. If $\phi^{*}$ has real rank zero, then it follows by Corollary 2.2 that $\phi^{*}$ factors through a commutative $\mathrm{C}^{*}$-algebra of real rank zero. This algebra has nuclear dimension equal to zero by [10, Proposition 1.1] and [53, Proposition 2.4] (see also [12, Theorem 2.6]). The fact that (iii) implies (i) is immediate. ∎
We end this section by applying Theorem 2.1 to obtain a class of $*{}^{*}$-homomorphisms with real rank zero.
@@
$\pi(g\cdot y)=g\cdot y+Y_{c}=y+Y_{c}=g\cdot(y+Y_{c})=g\cdot\pi(y).$
@@
$\psi(g\cdot(y+Y_{c}))=\psi(y+Y_{c})=\phi(y)=\phi(g\cdot y)=g\cdot\phi(y)=g\cdot\psi(y+Y_{c}).$
Hence, the $*{}^{*}$-homomorphism $\phi^{*}$ factors equivariantly as
@@
[…]
Since $G$ acts trivially on $Y_{c}$, we have that $C(Y_{c})\rtimes_{r}G=C(Y_{c})\otimes\mathrm{C}^{*}_{r}(G)$. Then, $G$ is compact connected, so $\mathrm{C}^{*}_{r}(G)$ has real rank zero by [31, Theorem 1]. Moreover, $C(Y_{c})$ has real rank zero by [10, Proposition 1.1]. In particular, $C(Y_{c})$ is an AF-algebra. Then, $C(Y_{c})\otimes\mathrm{C}^{*}_{r}(G)$ has real rank zero by [10, Theorem 3.2] and we reach the conclusion. ∎
@@ 3. Permanence properties
By assumption, $A_{n}\subseteq B_{n}$ has real rank zero, so there exists an invertible self-adjoint $b_{n}\in B_{n}$ such that $\|a_{n}-b_{n}\|\leq\epsilon/2$. Since $\psi_{n,\infty}$ is a unital $*{}^{*}$-homomorphism, $\psi_{n,\infty}(b_{n})$ is a self-adjoint invertible element. Then, triangle inequality gives that $\|a-\psi_{n,\infty}(b_{n})\|\leq\epsilon$. Hence $A\subseteq B$ has real rank zero by Theorem 1.6. ∎
@@
We will now examine the behaviour of inclusions of real rank zero under extensions. The next three propositions follow the proof of [10, Theorem 3.14].
@@
Let $x\in A_{+}$ be nonzero. By Corollary 3.3, the inclusion $\overline{xAx}\subseteq\overline{xBx}$ has real rank zero. Therefore, replacing the inclusion $A\subseteq B$ with $\overline{xAx}\subseteq\overline{xBx}$, it suffices to check that any nonzero projection in $\theta(A/I)$ lifts to a projection in $B$. Let $a\in A$ be a self-adjoint element of norm $1$ such that $a+J$ is a nonzero projection and let $10<\epsilon<$ 1. Then there exists $b\in B$ self-adjoint with finite spectrum and norm $1$ such that $\|a-b\|\leq\epsilon^{2}$. Since $b$ is a finite linear combination of its spectral projections, we can write $b=\sum\limits_{k=1}^{n}\lambda_{k}p_{k}$, for some $n\in\mathbb{N}$, $\lambda_{k}\in\mathbb{R}$, and $p_{k}\in B$ pairwise orthogonal projections.
@@
(3.1) $\Big{\|}\pi\Big{(}\sum(\lambda_{k}-\lambda_{k}^{2})p_{k}\Big{)}\Big{\|}=\|\pi(b-b^{2}-a+a^{2})\|\leq 3\epsilon^{2}.$
@@
Then, $\|\pi(p-a)\|\leq 2\epsilon+\epsilon^{2}$, so there is a positive contraction $e\in J$ such that $\|(1-e)(p-a)\|\leq 3\epsilon+\epsilon^{2}$. Hence, $z=(1-e)a+ep\in B$ is self-adjoint such that $z-a\in J$ and $\|p-z\|\leq 3\epsilon+\epsilon^{2}$. So for $\epsilon$ small enough we have $\|p-z\|<\frac{1}{16}$. Since $\|z+p\|\leq 2+\frac{1}{16}$, it follows that
@@
Let $x,y\in A^{\dagger}$ be positive orthogonal elements and $\epsilon>0$. Since $xy=0$, at least one of them is in $A$, so let us assume that $x\in A$. Let $C=\overline{xAx}$ and $\pi:A\to A/I$ be the canonical quotient map. Then $\pi(C)$ is the hereditary subalgebra of $A/I$ generated by $\pi(x)$. By applying Corollary 3.3 to the inclusion $A/I\subseteq A/I$ and the element $\pi(x)$, it follows that $\pi(C)$ has real rank zero. Therefore, there exists a projection $q\in\pi(C)$ such that
@@
By assumption, $q$ lifts to a projection $p$ in $B$, which can be assumed to be in $\overline{xBx}$ by [10, Lemma 3.13].
@@ 4. K-theoretic properties
In this section we characterise inclusions of real rank zero by appealing to the pairing between $K$-theory and traces. The motivation for this approach comes from [43], where Rørdam proved that for a certain class of $\mathrm{C}^{*}$-algebras, real rank zero is equivalent to the image of the pairing map in $K$-theory to be uniformly dense in the set of affine functions on the trace space ([43, Theorem 7.2]). The proof of the next result essentially follows the strategy in [43, Proposition 7.1] and [42, Theorem 7.2].
@@
where $\tau_{n}$ is the non-normalized extension of $\tau$ to $M_{n}(C)$. Following [7, Definition 1.5], we say that $C$ has strict comparison of positive elements with respect to tracial states if whenever $a,b\in M_{n}(C)_{+}$ for some $n\in\mathbb{N}$ and $a\in\overline{M_{n}(C)bM_{n}(C)}$, we have
@@
for all $\tau\in T(B)$. Since $B$ has strict comparison of positive elements with respect to tracial states, we get that $g=[q]_{0}\in K_{0}(B)_{+}$, for some projection $q\in M_{n}(B)$. Moreover, since $d_{\tau}(q)=\tau(q)$, strict comparison of $B$ yields that
@@
Furthermore, $f_{\delta/8}(a)$ is full in $B$, so the hereditary subalgebra $\overline{f_{\delta/8}(a)Bf_{\delta/8}(a)}$ is stably isomorphic to $B$ by [9, Theorem 2.8]. Since stable rank one is preserved by stable isomorphisms ([41, Theorem 3.6]), it follows that $\overline{f_{\delta/8}(a)Bf_{\delta/8}(a)}$ has stable rank one. Then, one can follow the last part of the proof of [42, Theorem $7.2$] to obtain a projection $p\in B$ such that $f_{\delta}(a)\leq p\leq f_{\delta/16}(a).$
Let $n\in\mathbb{N}$ and consider the unital inclusion $M_{n}(A)\subseteq M_{n}(B)$. Since $A\subseteq B$ is full, so is the inclusion $M_{n}(A)\subseteq M_{n}(B)$ ([23, Lemma 3.16]). Moreover, strict comparison passes to matrix amplifications, and so does stable rank one ([41, Theorem 3.3]). Lastly, one also has that $\gamma^{*}\left(\mathrm{Aff}(T(M_{n}(A)))\right)\subseteq\overline{\rho_{M_{n}(B)}(K_{0}(M_{n}(B)))}$. Hence the proof above shows that the inclusion $M_{n}(A)\subseteq M_{n}(B)$ has real rank zero. Thus, $A\otimes\mathcal{K}\subseteq B\otimes\mathcal{K}$ has real rank zero by Proposition 3.1. ∎
@@
If $T(B)$ is the empty set, then the conclusion is vacuosly true, so let us assume that $T(B)\neq\emptyset$. Let $f$ be a strictly positive continuous affine function on $T(A)$. Moreover, $A$ is unital and exact, so all quasitraces are traces by [27]. Then, there exists a positive element $a\in A\otimes\mathcal{K}$ such that $d_{\tau}(a)=f(\tau)$ for all $\tau\in T(A)$ ([2, Theorem 7.14] see also the second theorem in [2, pp.38]). Since the inclusion $A\otimes\mathcal{K}\subseteq B\otimes\mathcal{K}$ has real rank zero, there exists an increasing approximate unit of projections $(p_{n})_{n\geq 1}$ for the hereditary subalgebra generated by $a\in A\otimes\mathcal{K}$ in $B\otimes\mathcal{K}$ ([55, Proposition 1.2]).
@@
In fact, further assuming strict comparison on $B$ we can strengthen the result in Theorem 4.2. Similarly to [43, Theorem 7.2], we can show that projections coming from $A\otimes\mathcal{K}$ can be weakly divided in $K_{0}(B)_{+}$.

### v3

@@ 2. Inclusions of commutative $\mathrm{C}^{*}$-algebras
The if direction follows from Lemma 1.11 and [11, Proposition 1.1]. Conversely, let us suppose that the inclusion has real rank zero. We first claim that $\phi$ is constant on each connected component of $Y$. To prove it, let $Y_{0}$ be a connected component of $Y$ and $y_{1},y_{2}\in Y_{0}$. If $Y_{0}$ is a singleton, we are done, so let us assume that $y_{1}$ and $y_{2}$ are distinct. Suppose for a contradiction that $\phi(y_{1})\neq\phi(y_{2})$. By Urysohn’s lemma, we then pick a self-adjoint $f\in C(X)$ such that $f(\phi(y_{1}))<0$ and $f(\phi(y_{2}))>0$. Let $\epsilon>0$ such that $2\epsilon<\min\{f(\phi(y_{2})),|f(\phi(y_{1}))|\}$.
Since $\phi^{*}$ has real rank zero, there exists an invertible self-adjoint $g\in C(Y)$ such that $|g(y)-\phi^{*}(f)(y)|\leq\epsilon$ for all $y\in Y$. In particular,
@@
commutes, where $\psi^{*}$ and $\pi^{*}$ are the dual maps of $\psi$ and $\pi$ respectively. Since $Y$ is compact and Hausdorff, it follows that $Y_{c}$ is compact and totally disconnected. It is well-known that $Y_{c}$ is Hausdorff. To see this, note that in any compact Hausdorff space, connected components are quasicomponents and hence closed (see for example [18, Proposition 1.4]). As $Y$ is compact and Hausdorff, and in particular normal, any two connected components are contained in disjoint open sets. Consequently, the quotient space $Y_{c}$ is Hausdorff. Hence, $C(Y_{c})$ has real rank zero by [11, Proposition 1.1]. Moreover, $\pi$ is surjective, which implies that $\pi^{*}$ is injective by duality. Hence $\pi^{*}(C(Y_{c}))$ has real rank zero. Since $C(X)\subseteq\pi^{*}(C(Y_{c}))\subseteq C(Y)$, we reach the conclusion. ∎
@@
Let $X,Y$ be compact Hausdorff spaces and $\phi:Y\to X$ be a continuous map which induces a -homomorphism $\phi^{*}:C(X)\to C(Y)$. Then $\phi^{*}$ has real rank zero if and only if it factors through a commutative $\mathrm{C}^{*}$-algebra of real rank zero.
@@
The if direction follows from Lemma 1.10. Conversely, suppose that $\phi^{*}$ has real rank zero. Note that $\phi^{*}(C(X))$ is a commutative $\mathrm{C}^{*}$-algebra, so we can identify it with some $\mathrm{C}^{*}$-algebra $C(Z)$ by Gelfand duality. Then the conclusion follows by applying Theorem 2.1 to the inclusion $C(Z)\subseteq C(Y)$ and [11, Proposition 1.1]. ∎
We use Corollary 2.2 to characterise -homomorphisms between commutative $\mathrm{C}^{*}$-algebras which have nuclear dimension equal to zero. Let us first record the following definition.
@@
([51, 9]) Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras and $\theta:A\to B$ be a -homomorphism. It is said that $\theta$ has nuclear dimension equal to zero if for any finite set $\mathcal{F}\subseteq A$ and $\epsilon>0$, there exist a finite dimensional $\mathrm{C}^{*}$-algebra $F$ and cpc maps
@@
Moreover, a $\mathrm{C}^{*}$-algebra $A$ is said to have nuclear dimension equal to zero if the identity map on $A$ has nuclear dimension equal to zero. It is known that a separable $\mathrm{C}^{*}$-algebra has nuclear dimension equal to zero if and only if it is approximately finite dimensional ([53, Theorem 4.2.3]).
@@
Let $A,B$ be unital $\mathrm{C}^{*}$-algebras and $\theta:A\to B$ be a unital -homomorphism with nuclear dimension equal to zero. Then $\theta$ has real rank zero.
@@
Let $a\in A$ be a self-adjoint contraction and $\epsilon>0$. Building on [53, Theorem 3.4], there exist a finite dimensional $\mathrm{C}^{*}$-algebra $F$, a ucp map $\psi:A\to F$, and a unital -homomorphism $\eta:F\to B$ such that $\|\theta(a)-\eta(\psi(a))\|<\epsilon/2$ ([14, Lemma 3.1]). [^6] Even though [14, Lemma 3.1] is only stated in the setting when $A$ and $B$ are separable, this assumption is not needed. Since $\psi(a)$ is self-adjoint and $F$ is finite dimensional, there exists a self-adjoint invertible $b^{\prime}\in F$ such that $\|\psi(a)-b^{\prime}\|\leq\epsilon/2$. Therefore, $\eta(b^{\prime})$ is self-adjoint invertible and
@@
Let $X,Y$ be compact Hausdorff spaces and $\phi:Y\to X$ be a continuous map which induces a -homomorphism $\phi^{*}:C(X)\to C(Y)$. The following are equivalent:
@@
The fact that (i) implies (ii) follows from Proposition 2.4. If $\phi^{*}$ has real rank zero, then it follows by Corollary 2.2 that $\phi^{*}$ factors through a commutative $\mathrm{C}^{*}$-algebra of real rank zero. This algebra has nuclear dimension equal to zero by [11, Proposition 1.1] and [55, Proposition 2.4] (see also [13, Theorem 2.6]). The fact that (iii) implies (i) is immediate. ∎
We end this section by applying Theorem 2.1 to obtain a class of -homomorphisms with real rank zero.
@@
$\pi(g\cdot y)=\pi(y)=g\cdot\pi(y).$
@@
$\psi(g\cdot\pi(y))=\psi(\pi(y))=\phi(y)=\phi(g\cdot y)=g\cdot\phi(y)=g\cdot\psi(\pi(y)).$
Hence, the -homomorphism $\phi^{*}$ factors equivariantly as
@@
$C(X)\rtimes_{r}G$$C(Y)\rtimes_{r}G$$C(Y_{c})\rtimes_{r}G.$$\scriptstyle{\tilde{\phi}}$
Since $G$ acts trivially on $Y_{c}$, we have that $C(Y_{c})\rtimes_{r}G\cong C(Y_{c})\otimes\mathrm{C}^{*}_{r}(G)$. Then, $G$ is compact connected, so $\mathrm{C}^{*}_{r}(G)$ has real rank zero by [33, Theorem 1]. Moreover, $C(Y_{c})$ has real rank zero by [11, Proposition 1.1]. In particular, $C(Y_{c})$ is an AF-algebra. Then, $C(Y_{c})\otimes\mathrm{C}^{*}_{r}(G)$ has real rank zero by [11, Theorem 3.2] and we reach the conclusion. ∎
@@ 3. Permanence properties
By assumption, $A_{n}\subseteq B_{n}$ has real rank zero, so there exists an invertible self-adjoint $b_{n}\in B_{n}$ such that $\|a_{n}-b_{n}\|\leq\epsilon/2$. Since $\psi_{n,\infty}$ is a unital -homomorphism, $\psi_{n,\infty}(b_{n})$ is a self-adjoint invertible element. Then, triangle inequality gives that $\|a-\psi_{n,\infty}(b_{n})\|\leq\epsilon$. Hence $A\subseteq B$ has real rank zero by Theorem 1.6. ∎
@@
We will now examine the behaviour of inclusions of real rank zero under extensions. The next three propositions follow the proof of [11, Theorem 3.14].
@@
Let $x\in A_{+}$ be nonzero. By Corollary 3.3, the inclusion $\overline{xAx}\subseteq\overline{xBx}$ has real rank zero. Therefore, replacing the inclusion $A\subseteq B$ with $\overline{xAx}\subseteq\overline{xBx}$, it suffices to check that any nonzero projection in $\theta(A/I)$ lifts to a projection in $B$. Let $a\in A$ be a self-adjoint element of norm $1$ such that $a+J$ is a nonzero projection and let $0<\epsilon<1$. Then there exists $b\in B$ self-adjoint with finite spectrum and norm $1$ such that $\|a-b\|\leq\epsilon^{2}$. Since $b$ is a finite linear combination of its spectral projections, we can write $b=\sum\limits_{k=1}^{n}\lambda_{k}p_{k}$, for some $n\in\mathbb{N}$, $\lambda_{k}\in\mathbb{R}$, and $p_{k}\in B$ pairwise orthogonal projections.
@@
(3.1) $\left\|\pi\left(\sum\limits_{k=1}^{n}(\lambda_{k}-\lambda_{k}^{2})p_{k}\right)\right\|=\|\pi(b-b^{2}-a+a^{2})\|\leq 3\epsilon^{2}.$
@@
Then, $\|\pi(p-a)\|\leq 2\epsilon+\epsilon^{2}$, so there is a positive contraction $e\in J$ such that $\|(1-e)(p-a)\|\leq 3\epsilon+\epsilon^{2}$. Hence, $z=(1-e)a(1-e)+ep+pe-epe\in B$ is self-adjoint such that $z-a\in J$ and $\|p-z\|\leq 3\epsilon+\epsilon^{2}$. So for $\epsilon$ small enough we have $\|p-z\|<\frac{1}{16}$. Since $\|z+p\|\leq 2+\frac{1}{16}$, it follows that
@@
Let $x,y\in A^{\dagger}$ be positive orthogonal elements and $\epsilon>0$. Since $xy=0$, at least one of them is in $A$, so let us assume that $x\in A$. Let $C=\overline{xAx}$ and $\pi:A\to A/I$ be the canonical quotient map. Then $\pi(C)$ is the hereditary subalgebra of $A/I$ generated by $\pi(x)$, so it has real rank zero. Therefore, there exists a projection $q\in\pi(C)$ such that
@@
By assumption, $q$ lifts to a projection $p$ in $B$, which can be assumed to be in $\overline{xBx}$ by [11, Lemma 3.13].
@@ 4. K-theoretic properties
In this section we characterise inclusions of real rank zero by appealing to the pairing between $K$-theory and traces. The motivation for this approach comes from [45], where Rørdam proved that for a certain class of $\mathrm{C}^{*}$-algebras, real rank zero is equivalent to the image of the pairing map in $K$-theory to be uniformly dense in the set of affine functions on the trace space ([45, Theorem 7.2]). The proof of the next result essentially follows the strategy in [45, Proposition 7.1] and [44, Theorem 7.2].
@@
where $\tau_{n}$ is the non-normalized extension of $\tau$ to $M_{n}(C)$. Following [8, Definition 1.5], we say that $C$ has strict comparison of positive elements with respect to tracial states if whenever $a,b\in M_{n}(C)_{+}$ for some $n\in\mathbb{N}$ and $a\in\overline{M_{n}(C)bM_{n}(C)}$, we have
@@
for all $\tau\in T(B)$. Since $B$ has strict comparison of positive elements with respect to tracial states, we get that $g=[q]_{0}\in K_{0}(B)$, for some projection $q\in M_{n}(B)$. [^7] This is a standard trick and goes as follows: write $g=[q^{\prime}]_{0}-[p^{\prime}]_{0}$ for projections $p^{\prime},q^{\prime}\in M_{n}(B)$ for some $n\in\mathbb{N}$. As $\tau(g)>0$ for all $\tau\in T(B)$ we have $\tau_{n}(p^{\prime})<\tau_{n}(q^{\prime})$ for all $\tau\in T(B)$. Hence $p^{\prime}$ is subequivalent to $q^{\prime}$ so $p^{\prime}\sim q^{\prime\prime}\leq q^{\prime}$ for some projection $q^{\prime\prime}$. Letting $q=q^{\prime}-q^{\prime\prime}$, we have $g=[q]_{0}\in K_{0}(B)$. Moreover, since $d_{\tau}(q)=\tau(q)$, strict comparison of $B$ yields that
@@
Furthermore, $f_{\delta/8}(a)$ is full in $B$, so the hereditary $\mathrm{C}^{*}$-subalgebra $\overline{f_{\delta/8}(a)Bf_{\delta/8}(a)}$ is stably isomorphic to $B$ by [10, Theorem 2.8]. Since stable rank one is preserved by stable isomorphisms ([43, Theorem 3.6]), it follows that $\overline{f_{\delta/8}(a)Bf_{\delta/8}(a)}$ has stable rank one. Then, one can follow the last part of the proof of [44, Theorem $7.2$] to obtain a projection $p\in B$ such that $f_{\delta}(a)\leq p\leq f_{\delta/16}(a).$
Let $n\in\mathbb{N}$ and consider the unital inclusion $M_{n}(A)\subseteq M_{n}(B)$. Since $A\subseteq B$ is full, so is the inclusion $M_{n}(A)\subseteq M_{n}(B)$ ([25, Lemma 3.16]). Moreover, strict comparison passes to matrix amplifications, and so does stable rank one ([43, Theorem 3.3]). Lastly, one also has that $\gamma^{*}\left(\mathrm{Aff}(T(M_{n}(A)))\right)\subseteq\overline{\rho_{M_{n}(B)}(K_{0}(M_{n}(B)))}$. Hence the proof above shows that the inclusion $M_{n}(A)\subseteq M_{n}(B)$ has real rank zero. Thus, $A\otimes\mathcal{K}\subseteq B\otimes\mathcal{K}$ has real rank zero by Proposition 3.1. ∎
@@
If $T(B)$ is the empty set, then the conclusion is vacuously true, so let us assume that $T(B)\neq\emptyset$. Let $f$ be a strictly positive continuous affine function on $T(A)$. Moreover, $A$ is unital and exact, so all quasitraces are traces by [29]. Then, there exists a positive element $a\in A\otimes\mathcal{K}$ such that $d_{\tau}(a)=f(\tau)$ for all $\tau\in T(A)$ ([2, Theorem 7.14] see also the second theorem in [2, pp.38]). Since the inclusion $A\otimes\mathcal{K}\subseteq B\otimes\mathcal{K}$ has real rank zero, there exists an increasing approximate unit of projections $(p_{n})_{n\geq 1}$ for the hereditary subalgebra generated by $a\in A\otimes\mathcal{K}$ in $B\otimes\mathcal{K}$ ([57, Proposition 1.2]).
@@
In fact, further assuming strict comparison on $B$ we can strengthen the result in Theorem 4.2. Similarly to [45, Theorem 7.2], we can show that projections coming from $A\otimes\mathcal{K}$ can be weakly divided in $K_{0}(B)_{+}$.